OpenStudy (anonymous):

PLEASE HELP ME ***need this to be 100% right please****

OpenStudy (anonymous):

@anx @Nurali

OpenStudy (michele_laino):

the inverse formula is: \[b _{1}=\frac{ 2A }{h }-b _{2}\] Please, try to substitute your measures and tell me your answer

OpenStudy (michele_laino):

no, please try another time

OpenStudy (michele_laino):

for example: \[b _{1}=\frac{ 2\times 91 }{ 7}-15=\]

OpenStudy (anonymous):

11

OpenStudy (michele_laino):

well done!

OpenStudy (anonymous):

thank you :)

OpenStudy (michele_laino):

thank you!